OpenStudy (umulas):

Find the exact values under the given condition cos(theta + beta) sin theta = 3/5, lies in quad I, sin beta = 5/13 lies in quad II Any idea on how to do this? Thanks

OpenStudy (iloveopenstudy2):



OpenStudy (iloveopenstudy2):

Now, can you find cos theta, and cos beta?

OpenStudy (iloveopenstudy2):

And then we use this formula: cos(α + β) = cos(α)cos(β) – sin(α)sin(β)

OpenStudy (umulas):

cos 4/5*-12/13 - 5/13*3/15 = -63/65 Thanks!

OpenStudy (iloveopenstudy2):

You're welcome!

OpenStudy (ebaar123):

cos(α + β) = cos(α)cos(β) – sin(α)sin(β)

OpenStudy (ebaar123):

so now just find the cos theta and cos beta